OpenStudy (anonymous):

n a certain store, the regular price of a refrigerator is 600 dollars. How much money is saved by buying this refrigerator at 20 percent off the regular price rather than buying it on sale at 10 percent off the regular price with an additional discount of 10 percent off the sale price? I got $6.00 by using the calculator but I want to know how to solve this step by step WITHOUT the calculator. Percentages confuse me...help please?

OpenStudy (shamil98):

600. 20% off = 480. 10% off reg price = 540 - 10% = 486. So yes, you are correct. 6 dollars are saved.

OpenStudy (jdoe0001):

well.. let's take a peek at the price with 10% off regular price MINUS 10% off FROM that discounted "sale" price so the fridge is 600 so \( \large \begin{array}{ccllll} amount&\% \\\hline\\ 600&100\\ x&10 \end{array}\implies \cfrac{600}{x}=\cfrac{100}{10}\implies \cfrac{60\cancel{ 00 }}{\cancel{ 100 }}=x\) so the discount is 60... so the "sale price" is 600-60 = 540 so how much is 10% off the "sale price" of 540? well \(\large \begin{array}{ccllll} amount&\% \\\hline\\ 540&100\\ x&10 \end{array}\implies \cfrac{540}{x}=\cfrac{100}{10}\implies \cfrac{54\cancel{ 00 }}{\cancel{ 100 }}=x\) so the discount is 54 more so 540 - 54 = \(\bf \color{red}{486}\) ------------------------------------------------------ so.. how much will it be by just taking 20% right off the 600 original price then? \(\large \begin{array}{ccllll} amount&\% \\\hline\\ 600&100\\ x&20 \end{array}\implies \cfrac{600}{x}=\cfrac{100}{20}\implies \cfrac{120\cancel{ 00 }}{\cancel{ 100 }}=x\) so is 120 so using 20% right off from the original price will be 600-120 = \(\bf \color{red}{480}\)

OpenStudy (anonymous):

Thank you both for explaining thoroughly ^-^ I just remembered the bat and ball technique lol

OpenStudy (jdoe0001):

yw